You are reading a book for your English class. It takes you 75 minutes to read 45 pages.
A) Find your average speed. Write an equation that relates the number of pages 'p' you read to the number of minutes 'm' you read.
B) The book you are reading has 195 pages. How long will it take you to read the book?

Answer by TimothyLamb(4379)

r = p/m
r = 45/75
r = 0.6 pages/minute
---
m = p/r
m = 195/0.6
m = 325 minutes
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---